50 ml N/2 HCl, 50ml N/2 H2SO4, and 200 ml N/10 NaOH are mixed. Calculate whether the mixture will be acidic or basic? What will be its normality?

Grade:

1 Answers

Vikas TU
4 years ago
Dear student
The total volume to be made is 1 litre which is 1000 litre hence now for each of the compounds the value of the mass*volume will be 50*1/10 = 5 millimoles for HCl,
For H2SO4 50*1/2 = 25 millimoles of H2So4 and 1/10*200 = 20 millimoles of NaOH so the total millimoles will be 20+25+5 = 50 of equivalent millimoles.
Now, the final mixture will be of 1000 litre and N normality so the  value of N will be 50/1000 = 1/20 = 0.05 normal of the solution will be formed.